The line 𝑦 = 𝑘𝑥 + 4, where k is a constant, is
graphed in the xy-plane. If the line contains the
point (c, d), where 𝑐 ≠ 0 𝑎𝑛𝑑 𝑑 ≠ 0, what is the slope
of the line in terms of c and d?

Answer by Alan3354(69443)

The line 𝑦 = 𝑘𝑥 + 4, where k is a constant, is
graphed in the xy-plane. If the line contains the
point (c,d), where 𝑐 ≠ 0 𝑎𝑛𝑑 𝑑 ≠ 0, what is the slope
of the line in terms of c and d?
------------------
At (c,d), d = k*c + 4
k = (d-4)/c
slope = (d-4)/c